OpenStudy (anonymous):

Find all polar coordinates of point P where P = (9,pi divided by 5). (9, pi divided by 5 + 2nπ) or (-9, pi divided by 5 + 2nπ) (9, pi divided by 5 + (2n + 1)π) or (-9, pi divided by 5 + 2nπ) (9, pi divided by 5 + nπ) or (-9, pi divided by 5 + nπ) (9, pi divided by 5 + 2nπ) or (-9, pi divided by 5 + (2n + 1)π)

OpenStudy (jdoe0001):

what they're really asking is all coterminal vectors that land at the same angle of \(\cfrac{\pi}{5}\)

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

so, which ones, well, just cycle through \(2\pi\) and you'll land there again, another \(2\pi\) and there you are, and once again and again and again